The equilibrium constant for the reaction :
H₂(g)+S(s)⇔H₂S(g) is 20.0 at 925K and 10.25 at 1000K, respectively.
To calculate the standard enthalpy of reaction.
1
Expert's answer
2019-12-03T08:26:40-0500
  1. ∂lnK / ∂T = ΔH/(RT2)
  2. After integration: ln(K2/K1) = ΔH/R • (1/T1-1/T2) = ΔH/R • (T2-T1)/(T1•T2)
  3. Hence ΔH equals: ΔH = ln(K2/K1) • R • (T1T2)/(T2-T1) = -0,668 • 8,314 • (9,25•105)/75 = - 68496 = -68,5 kJ/mol

Good luck!